# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v3

@@ A.3.3 Proof of Local Statistical Error Bound in Theorem 1
$\frac{12}{(\mu+\lambda)^{2}}\leq\frac{96\rho^{2}/(\mu^{2}N)+99\kappa^{2}R^{2}}{128\rho^{2}/N+132L^{2}R^{2}+2\rho^{2}/n}$ (73)
we have
$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}$ (74)
@@
$\lambda\geq\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu$ (75)
Combining two-case arguments, we have:
@@
$C_{3}$ $=\frac{192\rho^{2}}{\mu^{2}},$ (78)
$C_{4}$ $=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee 198\kappa^{2},$ (79)
@@
Combining the results in A.3.2 and A.3.3, we leave a remark below to further interpret the effect of personalization in obj (1) and discuss how to incorporate the conditions of $\lambda$ when establishing the global and local statistical accuracy.
Remark 2.
To better understand how such an interpolation is achieved through an adaptive personalization degree, we examine the role of $\lambda$ in navigating between the two extreme cases: GlobalTrain and LocalTrain. As the statistical heterogeneity $R\to\infty$, we have $\lambda\to 0$, leading to a high degree of personalization. The statistical error bound becomes $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, $\lambda$ increases, leading to a low degree of personalization. The statistical error bound eventually converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain under a homogeneous setting. As $R\in(0,\infty)$, $\lambda$ transits between the two extremes, making FedProx perform no worse than LocalTrain while benefiting from global training when clients share similarities.
@@
Appendix B Optimization Convergence Rate
@@ B.1 Algorithm
      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server;
      $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$;
Algorithm 1 One-stage FedProx
@@ B.2.1 Proof of Lemma 2
First, from mishchenko2023convergence, we can directly obtain the analytic properties of $F_{i}$. Next, the analytic properties of $h_{i}$ are direct results from Assumption 3 and 2. Finally, to analyze the property of the mapping $w_{\star}^{(i)}$, we just need to use the result from the above claim as we know the definition of ${\bm{w}}_{\star}^{(i)}(\cdot)$ in (9).
@@ B.2.3 Outer Loop convergence
Suppose Assumption 3 and 2 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
@@ B.3 Proof of Theorem 2
$\bullet$ Convergence Rate For the connection term in (103), if we set
@@
B.4 Proof of Corollary 1
where we have used the fact that
$K\geq 2+\frac{\lambda+L}{\lambda+\mu}\cdot\log(1056\kappa^{2}).$ (118)
B.5 Proof of The Local Model Convergence Rate in Theorem 2
@@
$\begin{split}\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}&\leq 2\|e_{t}^{(i)}\|^{2}+2\|\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})-\operatorname{prox}_{L_{i}/\lambda}(\widetilde{{\bm{w}}}^{(g)})\|^{2}\\ &\leq 2\|e_{t}^{(i)}\|^{2}+2L_{w}^{2}\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2},\end{split}$ (123)

### v4

@@ A.4.2 The Statistical Convergence of the Local Model with a Choice of $\lambda$
$\frac{12}{(\mu+\lambda)^{2}}\leq\frac{96\rho^{2}/(\mu^{2}N)+99\kappa^{2}R^{2}}{128\rho^{2}/N+132L^{2}R^{2}+2\rho^{2}/n},$ (86)
then we will have
$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}.$ (87)
@@
$\lambda\geq\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu.$ (88)
Combining two-case arguments, we have
@@
$\begin{split}C_{3}&=\frac{192\rho^{2}}{\mu^{2}},\\ C_{4}&=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee 198\kappa^{2},\\ \end{split}$ (91)
@@
Therefore, combining results from Section A.3.1 and A.3.2 we can obtain the upper bound in (9).
@@
To sum up, we give the one-line rate in (90) for Corollary 1 and provide the potential solution of $\lambda$ with a closed-form solution in (92).
A.5 Discussion on the Statistical Convergence Rate and Comparison with Existing Results
Combining the results in A.3.1 and A.3.2, we leave a remark below to further interpret the effect of personalization in Problem (2) and discuss how to incorporate the conditions of $\lambda$ when establishing the global and local statistical accuracy.
To better understand how such an interpolation is achieved through an adaptive personalization degree, we examine the role of $\lambda$ in navigating between the two extreme cases: GlobalTrain and LocalTrain. As the statistical heterogeneity $R\to\infty$, we have $\lambda\to 0$, leading to a high degree of personalization. The statistical error bound becomes $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, $\lambda$ increases, leading to a low degree of personalization. The statistical error bound eventually converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain under a homogeneous setting. As $R\in(0,\infty)$, $\lambda$ transits between the two extremes, making PFL Problem (2) perform no worse than LocalTrain while benefiting from global training when clients share similarities.
Our contribution can be summarized into three aspects :
$\bullet$ Our results are derived under more realistic assumptions. Chen et al. (2023c) imposes additional stringent conditions including a bounded parameter space $D$ and a uniform bound on the loss function $\ell(\cdot,z)$. Such conditions are difficult to satisfy in practice. For example, even a simple linear regression model with sub-Gaussian covariates would violate these assumptions.
$\bullet$ Our result is established for both global and local models with the same choice of $\lambda$. In contrast, Chen et al. (2023c) proposes a two-stage process where the global model is first estimated, followed by an additional local training phase with a different choice of $\lambda$. This two-step approach is likely the artifact of their analysis, as existing work on Problem (2) also shows that a single-stage optimization with a chosen $\lambda$ suffices to achieve the desired performance (T Dinh et al., 2020).
$\bullet$ We establish a state-of-the-art minimax statistical rate. As shown in Table 3, the established upper bound matches the lower bound established in (Chen et al., 2023c). In a heterogeneous case when $R\geq n^{-1/2}$, Chen et al. (2023c) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$. This bound becomes increasingly loose with larger $R$. In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned Problem 2 is always no worse than LocalTrain, independent of $R$. In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data. As the client number could be extremely large ($10^{5}$ devices in (Chen et al., 2023a)), order of $m$ is non-trivial. During the transition period when $m^{-1}n^{-1/2}<R<n^{-1/2}$, we achieve a rate of $\mathcal{O}(R^{2})$, again matching the lower bound, and is strictly faster than the rate $\mathcal{O}(n^{-1/2}R)$ established in Chen et al. (2023c).
Table 3: Results Comparison. $D:=\sup_{\bm{w},\bm{w}^{\prime}\in\mathcal{D}}\|\bm{w}-\bm{w}^{\prime}\|$ is the diameter of the parameter space ($D\geq R$) and $\|\ell\|_{\infty}=\inf\{M\in\mathbb{R}:\ell(\cdot,z)\leq M,\text{for any }z\}$ is a uniform upper bound on the loss function. Statistical rate refers to $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, with constants neglected.
Source Assumption Paradigm Statistical Rate
(Chen et al., 2023c) A1,2,3, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
Ours A1,2,3 One-stage $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
Lower Bound - - $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
Next, we discuss our theoretical contributions. Prior work (Chen et al., 2023c) established the rate through algorithmic stability, and to obtain bounded stability, the boundedness of the loss function is needed. This assumption (or bounded gradients) is, in fact, commonly imposed for analysis based on the tool of algorithmic stability. To relax this assumption and obtain a bound independent of the diameter of the domain/gradient norm/loss function norm, we need to jump out of the stability analysis framework and develop new techniques.
Our analysis does not rely on any algorithm but directly tackles the objective function, establishing rates using purely the properties of the loss. Specifically, using the strong convexity of the loss function, we first proved the estimation error is bounded by the gradient variance and an extra statistical heterogeneity term controlled by $\lambda$, as detailed in Lemma 1. This allows us to show for large $R$, a small $\lambda$ can be chosen to yield rate $O(1/n)$ and match one of the worst cases in the lower bound. For the complementary case $R\leq 1/\sqrt{n}$, we leveraged the GlobalTrain solution $\widetilde{\bm{w}}_{GT}$ as a bridge and proved the solutions of Problem (2) to $\widetilde{\bm{w}}_{GT}$ are bounded by a term inversely proportional to $\lambda$ (cf. Lemma 4), implying that they can be made arbitrarily close to $\widetilde{\bm{w}}_{GT}$ by setting $\lambda$ small. This, together with the rate of $\widetilde{\bm{w}}_{GT}$ as proved in Lemma 5, yields a rate of $O(1/N+R^{2})$. Combining the two cases, we proved minimax optimality.
Appendix B Algorithmic Convergence Rate
In this section, we will start to prove the algorithmic convergence rate rigorously. First, we present our Algorithm 1 in Appendix 1. In Appendix B.2, we provide some useful lemmas to facilitate our analysis, including the proof of Lemma 1 presented in the main paper and the inner loop and outer loop error contractions in Algorithm 1. Next, in Appendix B.3, we are ready to prove Theorem 2 with the local and global model convergence. In Appendix B.4, to show the communication cost and computation cost of our method, the proof of Corollary 2 is established. Finally, in Appendix B.5, as stochastic gradient descent would reduce the computation cost over the full sample, we present Algorithm 2 for the noise setting and provably show the benefit of collaboration to reduce the noise.
@@ B.1 Algorithm
Algorithm 1 Federated Gradient Descent with K-Step Local Optimization
@@
      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server  $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$
@@ B.2.1 Proof of Lemma 1
First, from (Mishchenko et al., 2023), we can directly obtain the analytic properties of $F_{i}$. Next, the analytic properties of $h_{i}$ are direct results from Assumption 2 and 1. Finally, to analyze the property of the mapping $w_{\star}^{(i)}$, we just need to use the result from the above claim as we know the definition of ${\bm{w}}_{\star}^{(i)}(\cdot)$ in (11).
@@ B.2.2 Inner Loop Convergence
Suppose Assumption 2 and 1 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
@@ B.2.3 Outer Loop convergence
Suppose Assumption 2 and 1 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
@@ B.3 Proof of Theorem 2
B.3.1 Proof of The Global Model Convergence Rate in Theorem 2
@@
$\bullet$ Convergence Rate. For the connection term in (114), if we set
@@
B.3.2 Proof of The Local Model Convergence Rate in Theorem 2
@@
$\begin{split}\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}&\leq 2\|e_{t}^{(i)}\|^{2}+2\|\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}_{t}^{(g)})-\operatorname{prox}_{L_{i}/\lambda}(\widetilde{{\bm{w}}}^{(g)})\|^{2}\\ &\leq 2\|e_{t}^{(i)}\|^{2}+2L_{w}^{2}\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2},\end{split}$ (130)